OpenStudy (anonymous):

Identify the vertex and axis of symmetry. y=2x^2+16x+31

OpenStudy (anonymous):

I thought you did -b/2a to find the x coordinate but I am confused as to if that is correct or not

OpenStudy (dumbcow):

that is correct x = -b/2a x = -16/4 = -4 substitute that in for x to get y y = 2(-4)^2 +16(-4) +31 y =-1 vertex: (-4,-1)